Thomas travels 3 miles to get to school. This is one-fourth the distance that tammy travels. How far does tammy travel
Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

Thomas travels 3 miles to get to school. This is one-fourth the distance that tammy travels. How far does tammy travel
----------
If 3 miles is 1/4, how many miles is 4/4?

Answer by addingup(3677)

3 miles equals what?
The problem says that 3 miles is 1/4 Tammy
How far does Tammy travel? We don't know, let's call it x:
3 = 1/4x
or 1/4x = 3
x = 3*4/1 (see note below)
x = 12 Tammy travels 12 miles
~~~~~~~~~~~~~~~~~~~~~~~~~
Note: 1/4 was multiplying on the left. We divide both sides by 1/4
(1/4x)/1/4 = 3/1/4 On the left 1/4 on the numerator and 1/4 on the denominator cance each other, and you are left with:
x = 3/(1/4)
So here is the twist: dividing by a fraction is the same as multiplying by its reciprocal. Take the easiest, 1/2. We all understand 1/2. If you have 10, 1/2 is 5, if you have 12, 1/2 is 6, etc. In decimals, 1/2 is 0.5. So, 0.5 x 10 = 5, 0.5 x 12 = 6. Now let's divide (use your calculator to follow along): 10/0.5 = 20 and 12/0.5 = 24 --ANY TIME YOU DIVIDE A NUMBER BY A FRACTION OR DECIMAL, YOU GET A BIGGER NUMBER-- Now try with the fraction: 10*2/1 = 20; 12*2/1 = 24. If you have questions, email me. I'll be glad to help.